You are to receive a $100,000 lump sum payment at the end of a 6-year period. The relevant interest rate or rate of return for the next six years is 3.00%, and the compounding period is semi-annually. What is the Present Value of the $100,000 lump sum payment.
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

100,000=P*(1+0.03/2)^(2*6)

P=100,000/(1+0.03/2)^(2*6)

=$100,000*0.836387421

=$83,638.74(Approx).